OpenStudy (anonymous):

A ceramics workshop makes wreaths, trees, and sleighs for sale at Christmas. A wreath takes 3 hours to prepare, 2 hours to paint, and 9 hours to fire. A tree takes 14 hours to prepare, 3 hours to paint, and 4 hours to fire. A sleigh takes 4 hours to prepare, 15 hours to paint, and 7 hours to fire. If the workshop has 116 hours for prep time, 64 hours for painting, and 110 hours for firing, How many of each can be made?

OpenStudy (anonymous):

A. 8 wreaths, 6 trees, 2 sleighs B. 6 wreaths, 2 trees, 8 sleighs C. 9 wreaths, 7 trees, 3 sleighs D. 2 wreaths, 8 trees, 6 sleighs

OpenStudy (anonymous):

call number of wreaths w , trees t and sleighs s. now you have 3 equations with 3 unknowns: 3w + 14t + 4s = 116 2w +3t + 15s = 64 9w+4t+7s=110

OpenStudy (anonymous):

i thought u cant add unlike terms

OpenStudy (anonymous):

each equation is for different "process"

OpenStudy (anonymous):

ya i have no clu w/ this one

OpenStudy (anonymous):

do you know how to solve these equations ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

another way is plugging the answers into the equations .. the answer that satisfies all 3 equations is the correct one

OpenStudy (anonymous):

ohh ok gimmie a min to figure it out .. sry this ones easy went right over my head

OpenStudy (anonymous):

but solving them is better since you wont always have answers. so lets do it using both elimination and substitution. if you multiply the first equation by 3 then subtract the 3rd one from the first you have: 38t+5s=238

OpenStudy (anonymous):

ok. we can keep on solving it if you wish

OpenStudy (anonymous):

it would b a correct?

OpenStudy (anonymous):

yes, A is the correct answer

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

but, it is really good practice to solve the equations by hand and not substituting answers

OpenStudy (anonymous):

but, it's up to you :)

OpenStudy (anonymous):

its hard for me on the computer ... hence y im having some trouble in this class but im almost done thank god